OpenStudy (anonymous):

Can someone please help me? The position of an object at time t is given by s(t) = 3 - 4t. Find the instantaneous velocity at t = 8 by finding the derivative.

OpenStudy (mrnood):

start with the equation s=3-4t can you find ds/dt ?

OpenStudy (anonymous):

What is ds?

OpenStudy (anonymous):

I know 3 is 0

OpenStudy (anonymous):

Can I do s=3-4t s=0-4t 0=-4t t=-4

OpenStudy (anonymous):

is a derivative always the coefficient of x?

OpenStudy (anonymous):

@Asevilla5

OpenStudy (asevilla5):

the derivative of a constant number (like 3) is 0 the derviative of a number times a t (like -4t) is just the number (-4) since the position function is a straight line, then the velocity will be constant for all time so there will always be a number -4 for a line you can calculate the slope by using two different points (rise over run) but a derivative is like moving those two points so close together its like taking the slope of a single point. since the slope is being taken at a single point, there is no ∆ x like a two point equation (rise/run = ∆y/∆x) so its the slope of a point at a single instant in time

OpenStudy (anonymous):

Thank you for that explanation, but i also want to know if the coefficient of x is always the derivative?

OpenStudy (anonymous):

velocity is constant, it will be always -4m/sec

OpenStudy (anonymous):

ohhh because it's the slope so it will always rise 4 and go left 1?

OpenStudy (asnaseer):

@lolabieber - you might find tis useful for understanding derivatives: http://www.mathsisfun.com/calculus/derivatives-introduction.html

OpenStudy (anonymous):

okay thank you guys for all your help :) I appreciate it!

OpenStudy (asevilla5):

@lolabieber yes